In numbers 5-7,

(a) make a table of values of t, x and y based on the specified intervals of the given parametic equations

(b) sketch the plane curve defined by the given parametic equations at the specified interval

(c) find dy/dx without eliminating the parameter t. simplify your answers

(d) find an equation such that the parameter t is eliminated


5. x=t²+1 and y=t³-2 for 0≤t≤2

6. x=sqrt(5t) and y=2t+2 for 0≤t≤5

7. x=cost and y=2-cos²t for 0≤t≤π


1
Expert's answer
2021-11-17T00:34:24-0500

5.

(a)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 t & x & y \\ \hline
 0 & 1 & -2 \\
 0.2 & 1.04 & -1.992 \\ 
 0.4 & 1.16 & -1.936 \\ 
 0.6 & 1.36 & -1.784 \\ 
 0.8 & 1.64 & -1.488 \\ 
 1 & 2 & -1 \\ 
 1.2 & 2.44 & -0.272 \\
 1.4 & 2.96 & 0.744\\ 
 1.6 & 3.56 & 2.096 \\ 
 1.8 & 4.24 & 3.832 \\ 
 2 & 5 & 6 \\ 
\end{array}"

(b)




(c)


"\dfrac{dy}{dx}=\dfrac{dy/dt}{dx/dt}=\dfrac{3t^2}{2t}=\dfrac{3}{2}t, t\not=0"

(d)


"t=\sqrt{x-1}, t\geq0"

"t=-\sqrt{x-1}, t<0"

"y(x)= \begin{cases}
 -(x-1)^{3/2}-2 &t<0 \\
 (x-1)^{3/2}-2 &t\geq0 
\end{cases}"

6.


(a)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 t & x & y \\ \hline
 0 & 0 & 2 \\
 0.5 & \sqrt{2.5} & 3 \\ 
 1 & \sqrt{5} & 4 \\ 
 1.5 & \sqrt{7.5} & 5 \\ 
 2 & \sqrt{10} & 6 \\ 
 2.5 & \sqrt{12.5} & 7 \\ 
 3 & \sqrt{15} & 8 \\
 3.5 & \sqrt{17.5} & 9\\ 
 4 & \sqrt{20} & 10 \\ 
 4.5 & \sqrt{22.5} & 11 \\ 
 5 & 5 & 12 \\ 
\end{array}"


(b)



(c)


"\dfrac{dy}{dx}=\dfrac{dy/dt}{dx/dt}=\dfrac{2}{\dfrac{5}{2\sqrt{t}}}=\dfrac{4}{5}\sqrt{t}, t>0"

(d)


"t=\dfrac{x^2}{5}, x\geq0"

"y(x)=\dfrac{2}{5}x^2+2, x\geq0"

7.

(a)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 t & x & y \\ \hline
 0 & 1 & 1 \\
 \pi/6 & \sqrt{3}/2 & 5/4 \\ 
 \pi/4 & \sqrt{2}/2 & 3/2 \\ 
 \pi/3 & 1/2 & 7/4 \\ 
 \pi/2 & 0 & 2 \\ 
 2\pi/3 & -1/2 & 7/4 \\ 
 3\pi/4 & -\sqrt{2}/2 & 3/2 \\
 5\pi/6 & -\sqrt{3}/2 & 5/4\\ 
 \pi & -1 & 1 \\ 
\end{array}"

(b)




(c)


"\dfrac{dy}{dx}=\dfrac{dy/dt}{dx/dt}=\dfrac{2\sin t\cos t}{-\sin t}=-2\cos t, t\not=\pi n, n\in \Z"

(d)

"y(x)=2-x^2, -1\leq x\leq1"